OpenStudy (kimo_indie):

What is the slope of the line that contains the points (2, -6) and (-3, 1)

OpenStudy (alexandervonhumboldt2):

use slope formula: (y1-y2)/(x1-x2)=substitute points=(-6-1)/(2+3)=(-7)/5=-7/5